Restore lab
[Paco, Tega, JC, Yehonathan]
We followed the instructions here. There were no major issues, apart from the fb1 ntp server sync taking long time after rebooting once.
ETMY damping
[Yehonathan, Paco]
We noticed that ETMY had to much RMS motion when the OpLevs were off. We played with it a bit and noticed two things: Cheby4 filter was on for SUS_POS and the limiter on ULCOIL was on at 0 limit. We turned both off.
We did some damping test and observed that the PIT and YAW motion were overdamped. We tune the gain of the filters in the following way:
SUSSIDE_GAIN 1250->50
SUSPOS_GAIN 200->150
SUSYAW_GAIN 60->30
These action seem to make things better.
|